Solve
2x³+6x²=20x
2x³+6x²-20x= 0
2x (x²+3x-10)
(x+5)(x-2)=0
I am unsure if the correct solution set is {-5, 0, 2}

Answer by ewatrrr(24785)

 
Hi,
2x³+6x²-20x= 2x (x²+3x-10)= 2x(x+5)(x-2)
2x(x+5)(x-2)=0
2x = 0 x = 0
0r (x+5)= 0 x = -5
0r (x-2)=0 x = 2
solution set {-5, 0, 2} is correct. Good work!